### v1

@@ 7.5 Hybrid Measurement
We define hybrid measurement as combining more than one of the aforementioned measurement methods. Currently, three existing CFA proposals fall into the category, which use a combination of hash-based measurement and abstract execution. LO-FAT [12] uses abstract execution for loop paths and hash-based measurement for the rest of the control-flow path; avoiding hash-based measurements for loops avoids the path explosion problem. OAT [6] is the first CFA scheme to explicitly address control-flow and data-only attacks on embedded systems. OAT addresses this through a combination of hash-based measurements and compact execution traces using software-only techniques in a TEE (ARM TrustZone).
@@
OAT’s system overview is given in Figure 18. The source program, $T_{\mathcal{P}}$, undergoes control-flow and critical variable instrumentation using an LLVM-based compiler. At run-time, a TEE-based measurement engine consumes application events through a trampoline library in the normal world. The engine checks the integrity of ‘critical variables’, which are identified by developers, using value-based approach to reduce performance overhead. More specifically, it checks whether the value at an instrumented load instruction is the same as the value recorded at its previous instrumented store instruction. The engine also measures CFI using a combination of traces for forward-edge control flow transfers (i.e. branches and indirect calls/jumps), and hash-based measurements for backward-edge transfers (returns) using BLAKE2. Traces and hashes are combined to form an ‘attestation blob’ that is sent to $\mathcal{V}$. After receiving the blob, $\mathcal{V}$ abstractly executes $T_{\mathcal{P}}$ using its CFG by starting from its entry BBL to reconstruct the measurement hatch, which we described previously in §7.4.
@@
ARI by Wang et al. [43] focusses on challenges in attesting autonomous cyber-physical systems. ARI relies on the decomposition of mission-critical software into sandboxed ‘compartments’, which are monitored using a combination of control- and data-flow techniques. Compartments are isolated using a memory protection unit or with software-based instrumentation. A trust anchor based on TEEs (ARM TrustZone or RISC-V Keystone) or a dedicated hardware module is suggested. ARI monitors control flows between compartments, e.g. hashes of return addresses on the stack; data flows of critical variables; transfers between compartments; and event timestamps. What is defined as a ‘critical’ component or data is set using a policy-based approach. These policies are translated into appropriate instrumentations using an LLVM-based compiler. During $T_{\mathcal{P}}$’s execution, a recording engine captures control and data flow events in the form of an event log, inspired by aircraft flight recorders. The workflow of ARI is given in Figure 19. In Figure 20, the verification procedure is given, whereby the application is re-executed and different control- and data-flow measurements are calculated and compared with taht sent by $\mathcal{P}$.
@@
    $\mathcal{P}$ sends a sequence of control-flow transfers and the final hash value as an attestation report upon request from $\mathcal{V}$.
@@
A second advantage is that it ameliorates a major drawback of hash-based measurements: the requirement for $\mathcal{V}$ to store all possible control-flow paths and their corresponding hash values. Return addresses, which serve as input of hash computation, are obtained by abstract execution based on memory addresses of function call. As such, $\mathcal{V}$ does not have to know learn possible backward control-flow paths during the offline phase, but instead only the final value.
@@ 8.2 Discussion and Analysis
Many CFA schemes claimed the ability to detect code injection attacks, i.e. (A.1). However, a significant minority of proposals (7/30, $\approx$23%) do not provide this due to features bestowed by the scheme itself. Rather, they make assumptions regarding on the availability of mechanisms that enforce some aspect of the target program’s immutability; for example, using executable-space, data execution prevention, and other memory protection mechanisms [12, 47]. LO-FAT, for instance, assumes that “the adversary cannot modify program code at run-time (marked as rx)” (p. 3, [12]) as an implicit requirement for a memory protection mechanism. Geden & Rasmussen [47] requires non-writable code regions and non-executable data regions. In a TEE-based scheme, OAT [6] assumes that the attackers cannot tamper with the instrumented code or the trampoline library in the normal world, citing some memory protection methods for embedded devices as potential solutions [93, 94]. [^7] We note that, in some cases, it is possible to bypass memory protection mechanisms, e.g. DEP, using ROP-style attacks. To this end, we refer the reader to work by Göktas et al. [95].
@@
In a related challenge, many proposals attest the program at launch-time, which can give rise to TOCTOU attacks if an adversary conducts an attack before or after an attestation challenge. Continuous reporting proposals explicitly aim to address this challenge [34, 44] (see §4.2), where $\mathcal{P}$ actively actively sends measurements to $\mathcal{V}$. However, this paradigm is used by a very small number of proposals (2/31, $\approx$6%). Conventional attestation schemes may still detect this by $\mathcal{P}$ storing measurements until they are requested by $\mathcal{V}$ (see ARCADIS [41]). Yet, this relies on $\mathcal{V}$ determining an appropriate timeframe to issue the challenge. Simply performing the challenge at launch-time, for example, would be insufficient in detecting transient attackers.
@@ 9.2 Limited Deployment Scenarios
During the course of this work, it became evident that CFA proposals tend to evaluated in laboratory conditions or using fixed datasets, e.g. SPEC CPU benchmarks, that do not replicate many real-world scenarios. In particular, we recognised that CFA schemes do not consider long-running applications, like those that control sensors and actuators. Such programs may execute a large number of control-flow transfers during run-time. For schemes that measure control-flow transfers, $\mathcal{V}$ might receive large attestation reports; in a path-based scheme, $\mathcal{V}$ may find it challenging, if not infeasible, to obtain all legitimate control-flow paths due to the enormous state space. We observed that very few proposals (e.g. OAT [6]) support customisable code attestation, which could alleviate this problem.
A large number of proposals rely on the submission of program inputs in order to trigger a measurement of traversed CFG notes. This paradigm may cause issues in applications where the program inputs do not traverse affected nodes, e.g. affected by a code injection attack. For example, consider a device that accepts a significant number of inputs, such as a robotic arm or another cyber-physical system ($\mathcal{P}$) that communicates with a remote control service ($\mathcal{V}$). What parameters should $\mathcal{V}$ submit to $\mathcal{P}$? Is it possible that the supplied parameters could unintentionally miss malicious code within a larger complex program? These questions, among others, have received limited attention in the literature.
@@ 9.5 Addressing Physical Attacks
Many papers deem physical attacks on the devices as outside the scope of their threat models [47, 8]. This is acceptable where we can safely assume that adversaries are strictly remote or software-based ones. However, many CFA schemes are designed for embedded systems, which may be deployed in largely unsupervised environments, such as industrial control systems and critical national infrastructure. The value of such devices to adversaries may bring increase the viability of physical attacks, such as fault injection and side-channel attacks. Indeed, the authors of OAT state: “We trust code running inside the Secure World (e.g., the measurement engine) and assume that attackers cannot break the TrustZone protection. We also trust our compiler and the trampoline code. We assume that attackers cannot inject code in the Normal World or tamper with the instrumented code or the trampoline library” [6]. It is widely known that TEEs are susceptible to hardware attacks [50]. We pose this as a significant gap in the literature, and urge the development of CFA schemes that consider such threats.
@@
    Research Question: To what extent can existing CFA schemes be made explicitly resilient to physical hardware attacks, such as fault injections?
@@ 9.6 Extended Attestation Paradigms
The vast majority of CFA schemes follow a traditional remote attestation with a single has a rich, longstanding history dating back to the inception of the Trusted Platform Modules (TPM) in the early-2000s. We recognise that some different approaches to attestation, which were developed for TPM-based platforms, have not yet been applied to the domain of control-flow attestation. In particular, property-based [61], direct anonymous [99], and mutual attestation [100] have not been fully adapted to CFA. We believe that these areas are fruitful avenues for researchers in the field.
@@ 10 Conclusion
In this paper, we presented a comprehensive analysis of control-flow attestation as it is currently realised in the state of the art, exploring the concepts, solutions, techniques and assumptions of 31 proposals in the field. We categorised existing CFA proposals based on their key features using a common set of criteria, including their instrumentation method, target control-flow transfers, attacks addressed, attestation paradigm, and more. Based on this, we presented a series of open challenges and recommendations for steering future research in the field. Through our work, we hope it can enhance the reader’s understanding of the capabilities and considerations involved in designing and analysing control-flow attestation schemes, and the challenges that lie ahead.

### v3

@@ 7.5 Hybrid Measurement
We define hybrid measurement as combining more than one of the measurement methods above. Currently, three existing CFA proposals fall into the category, which combine hash-based measurement and abstract execution. LO-FAT [12] uses abstract execution for loop paths and hash-based measurement for the rest of the control-flow path; avoiding hash-based measurements for loops avoids the path explosion problem. OAT [6] is the first CFA scheme to explicitly address control-flow and data-only attacks on embedded systems. OAT addresses this through a combination of hash-based measurements and compact execution traces using software-only techniques in a TEE (ARM TrustZone).
@@
OAT’s system overview is given in Figure 18. The source program, $T_{\mathcal{P}}$, undergoes control-flow and critical variable instrumentation using an LLVM-based compiler. At run-time, a TEE-based measurement engine consumes application events using a trampoline in the normal world. The engine checks the integrity of ‘critical variables’, which are annotated by developers, using a value-based approach to reduce performance overhead. More specifically, it checks whether the value at an instrumented load instruction is the same as that recorded at the previous store instruction. The engine also measures CFI using a combination of traces for forward-edge control flow transfers (i.e. branches and indirect calls/jumps), and hash-based measurements for backward-edge transfers (returns). Traces and hashes are combined to form an ‘attestation blob’ that is sent to $\mathcal{V}$. After receiving the blob, $\mathcal{V}$ abstractly executes $T_{\mathcal{P}}$ using its CFG by starting from its entry BBL to reconstruct the measurement hatch, which we described previously in §7.4.
@@
ARI by Wang et al. [43] focusses on attesting autonomous cyber-physical systems by decomposing mission-critical software into sandboxed ‘compartments’, which are monitored using a combination of control- and data-flow techniques. Compartments are isolated using a memory protection unit or with software-based instrumentation. A trust anchor based on TEEs (ARM TrustZone or RISC-V Keystone) or a dedicated hardware module is suggested. ARI monitors control flows between compartments, e.g. hashes of return addresses on the stack; data flows of critical variables; transfers between compartments; and event timestamps. What is defined as a ‘critical’ component or data is set using a policy-based approach. These policies are translated into appropriate instrumentations using an LLVM-based compiler. During $T_{\mathcal{P}}$’s execution, a recording engine captures control and data flow events in the form of an event log, inspired by aircraft flight recorders. The workflow of ARI is given in Figure 19. In Figure 20, the verification procedure is given, whereby the application is re-executed and different control- and data-flow measurements are calculated and compared with those sent by $\mathcal{P}$.
@@
    $\mathcal{P}$ sends a sequence of control-flow transfers and the final hash value as an attestation report to $\mathcal{V}$.
@@
A second advantage is that it ameliorates a major drawback of hash-based measurements: the requirement for $\mathcal{V}$ to store all possible control-flow paths and their corresponding hash values. Return addresses, which serve as input of hash computation, are obtained by abstract execution based on memory addresses of function calls. As such, $\mathcal{V}$ does not have to know possible backward control-flow paths during the offline phase, but instead only the final value.
@@ 8.2 Discussion and Analysis
Many CFA schemes claimed the ability to detect code injection attacks, i.e. (A.1). However, a significant minority of proposals (7/30, $\approx$23%) do not provide this due to features bestowed by the scheme itself. Rather, they make assumptions regarding the availability of mechanisms that enforce some aspect of the target program’s immutability; for example, using executable-space, data execution prevention, and other memory protection mechanisms [12, 47]. LO-FAT, for instance, assumes that “the adversary cannot modify program code at run-time (marked as rx)” (p. 3, [12]) as an implicit requirement for a memory protection mechanism. Geden & Rasmussen [47] requires non-writable code regions and non-executable data regions. In a TEE-based scheme, OAT [6] assumes that the attackers cannot tamper with the instrumented code or the trampoline library in the normal world, citing some memory protection methods for embedded devices as potential solutions [92, 93]. [^7] We note that, in some cases, it is possible to bypass memory protection mechanisms, e.g. DEP, using ROP-style attacks. To this end, we refer the reader to work by Göktas et al. [94].
@@
In a related challenge, many proposals attest the program at launch-time, which can give rise to TOCTOU attacks if an adversary conducts an attack before or after an attestation challenge. Continuous reporting proposals explicitly aim to address this challenge [34, 44] (see §4.2), where $\mathcal{P}$ actively sends measurements to $\mathcal{V}$. However, this paradigm is used by a very small number of proposals (2/31, $\approx$6%). Conventional attestation schemes may still detect this by $\mathcal{P}$ storing measurements until they are requested by $\mathcal{V}$ (see ARCADIS [41]). Yet, this relies on $\mathcal{V}$ determining an appropriate timeframe to issue the challenge. Simply performing the challenge at launch-time, for example, would be insufficient in detecting transient attackers.
@@ 9.2 Limited Deployment Scenarios
During the course of this work, it became evident that CFA proposals tend to be evaluated in laboratory conditions or using fixed datasets, e.g. SPEC CPU benchmarks, that do not replicate many real-world scenarios. In particular, we recognised that CFA schemes do not consider long-running applications, like those that control sensors and actuators. Such programs may execute a large number of control-flow transfers during run-time. For schemes that measure control-flow transfers, $\mathcal{V}$ might receive large attestation reports; in a path-based scheme, $\mathcal{V}$ may find it challenging, if not infeasible, to obtain all legitimate control-flow paths due to the enormous state space. We observed that very few proposals (e.g. OAT [6]) support customisable code attestation, which could alleviate this problem.
A large number of proposals rely on the submission of program inputs in order to trigger a measurement of traversed CFG nodes. This paradigm may cause issues in applications where the program inputs do not traverse affected nodes, e.g. affected by a code injection attack. For example, consider a device that accepts a significant number of inputs, such as a robotic arm or another cyber-physical system ($\mathcal{P}$) that communicates with a remote control service ($\mathcal{V}$). What parameters should $\mathcal{V}$ submit to $\mathcal{P}$? Is it possible that the supplied parameters could unintentionally miss malicious code within a larger complex program? These questions, among others, have received limited attention in the literature.
@@ 9.5 Addressing Physical Attacks
It is widely common for authors to deem physical attacks as outside the scope of their threat models, implicitly or explicitly. This is acceptable where we can safely assume that adversaries are strictly remote or software-based ones. However, many CFA schemes are designed for embedded systems, which may be deployed in largely unsupervised environments, such as industrial control systems and critical national infrastructure. The value of such devices to adversaries may increase the viability of physical attacks, such as fault injection and side-channel attacks. Indeed, the authors of OAT state: “We trust code running inside the Secure World (e.g., the measurement engine) and assume that attackers cannot break the TrustZone protection. We also trust our compiler and the trampoline code. We assume that attackers cannot inject code in the Normal World or tamper with the instrumented code or the trampoline library” [6]. It is widely known that TEEs are susceptible to hardware attacks [51]. We pose this as a significant gap in the literature, and urge the development of CFA schemes that consider such threats.
@@
    Research Question: To what extent can CFA schemes be made explicitly resilient to hardware attacks, such as fault injections and side-channel analysis?
@@ 9.6 Extended Attestation Paradigms
The vast majority of CFA schemes follow a traditional remote attestation paradigm using a single proving device and verification authority. We recognise that some different approaches to attestation may be of interest to researchers, which were developed for TPM-based platforms but have not been fully realised in the area of control-flow attestation. In particular, property-based [61], direct anonymous [98], and mutual attestation [99] have not been adopted as potential attestation models for CFA. We believe that these areas are fruitful avenues for researchers in the field.
@@ 10 Conclusion
In this paper, we presented a comprehensive analysis of control-flow attestation as it is currently realised in the state of the art, exploring the concepts, solutions, techniques and assumptions of 31 proposals in the field. We categorised existing CFA proposals based on their key features using a common set of criteria, including their instrumentation method, target control-flow transfers, attacks addressed, attestation paradigm, and more. Based on this, we presented a series of open challenges and recommendations for steering future research in the field. Through this work, we hope to enhance the reader’s understanding of the capabilities and considerations involved in designing and analysing control-flow attestation schemes, and the challenges that lie ahead.